OpenStudy (anonymous):

how do you simplify from ((7/(t-9))-(7/(x-9)))/(t-x) to get (-7/((t-9)(x-9)) ? What is the step that i'm not seeing between these two?

OpenStudy (anonymous):

is it \[\huge \frac{\frac{7}{t-9}-\frac{7}{x-9}}{t-x}\]

OpenStudy (anonymous):

yes thats the starting fraction

OpenStudy (anonymous):

\[\huge \frac{\frac{(7x-63)-(7t-63)}{(t-9)(x-9)}}{t-x}\] \[\huge \frac{\frac{7x-7t}{(t-9)(x-9)}}{t-x}\] \[\huge \frac{\frac{-7(t-x)}{(t-9)(x-9)}}{t-x}\]then (t-x) cancels, k?

OpenStudy (anonymous):

where does the 63 come from?

OpenStudy (anonymous):

7(x-9) and 7 (t-9) when i made denominators common (x-9)(t-9)

OpenStudy (anonymous):

ohh!!! ok thank you!!!